The perimeter of a rectangle is 32 inches. The area of the rectangle is 48 square inches. Find the dimension of the rectangle.
Answer by stanbon(75887)

The perimeter of a rectangle is 32 inches. The area of the rectangle is 48 square inches. Find the dimension of the rectangle.
------------------------
Equations:
Area: lw = 48
Per:: 32 = 2(l + w)
-----------------------
lw = 48
16 = l+w
----
w = 16-l
---
Substitute for "w" and solve for "l":
l(16-l) = 48
16l - l^2 = 48
l^2 - 16l + 48 = 0
-------------
l^2 -4l-12l + 48 = 0
l(l-4)-12(l-4) = 0
(l-4)(l-12) = 0
If length = 4 inches; then width = 12 inches
If length = 12 inches; then width = 4 inches
================================================
Cheers,
Stan H.
=============